How many three-digit odd numbers can be formed using the numbers 1,2,3,4,5,6, and 7?
Answer by sudhanshu_kmr(1152)


no. of ways to select a digit for unit place = 3 ( i.e - - X)
no. of ways to select a digit for second place = 5 (except that is on unit )
similarly no. of ways to select for third place = 4

total no. of odd numbers = 3 * 5 * 4 = 60